# Towards Any-Quality Image Segmentation via Generative and Adaptive Latent Space Enhancement

### v1

@@ 1 Introduction
Accurate visual perception  [maskrcnn, panoptic, yolact, videoseg, salient-detection, rssformer, glesam] that can operate under diverse and unpredictable real-world conditions remains a fundamental task in computer vision research. Robust performance is especially critical in various high-level visual applications, such as robotics, autonomous driving, and medical image analysis. The recently developed Segment Anything Models (SAMs), including SAM [sam] and SAM2 [sam2], serving as a foundational segmentation model, have gained significant influence within the community  [rsprompter, sam-med, sam-med2, med-sam2, sa2va] due to their outstanding zero-shot segmentation abilities. It can interactively segment any object in an image using visual prompts such as points and bounding boxes. SAM’s robust generalizability has led to breakthroughs and new paradigms in various downstream tasks, including remote sensing [rsprompter, sam-rs2, sam-rs3], automatic data annotation [samrs, sam-label, p2p], and medical image segmentation [sam-med, sam-med2].
@@
Based on our previous work [glesam], this paper presents a substantial extension with the following key contributions. (1) First, we propose GleSAM++, an enhanced version with the key principle of decoupling the reconstruction process into degradation-level prediction and degradation-aware reconstruction. This design addresses a key limitation of the previous model, thereby further improving performance across any-quality images. (2) Second, we introduce a degradation-aware adaptive enhancement mechanism that dynamically regulates the reconstruction intensity in the latent space based on degradation perception, enabling a more targeted enhancement process. (3) Furthermore, we conduct a more extensive experimental evaluation across a broader range of aspects and scenarios, including robustness under diverse prompts and performance in real-world settings, thereby providing a more systematic validation of the model’s effectiveness. (4) We also provide additional analysis of the underlying factors contributing to GleSAM++’s effectiveness, along with detailed algorithmic descriptions of both the training and inference processes. (5) Finally, we present comprehensive ablation studies and visual explanations, offering deeper insights into the mechanisms behind GleSAM++.
@@ 2.1 Segmentation on Low-Quality Images
Executing robust segmentation across various scenarios is a critical issue. Numerous studies [benchmark-robustness, improve-robust, sam-robustness1, robustsam] have highlighted significant performance degradation in conventional segmentation models and foundational SAMs when confronted with low-quality images with degradation. Many related studies [improve-robust, benchmark-robustness, lwfa, fifo, dense-gram] have been proposed to enhance the robustness of segmentation models against low-quality data. For instance, URIE [urie] enhances segmentation robustness under multiple image impairments by introducing classification-based constraints. QualNet [QualNet] achieves quality-agnostic feature extraction through a reversible encoding scheme. Meanwhile, FIFO [fifo] promotes the learning of fog-resilient features in segmentation frameworks via a fog-pass filtering mechanism. These methods primarily consider a single type of degradation. Recently, RobustSAM [robustsam] is introduced to enhance the robustness of the SAM against multiple image degradations through anti-degradation feature learning. However, its performance also struggles when dealing with complex degradations. The real-world image noise is often too complex to be modeled by a single degradation [image-noise, image-noise2, image-noise3]. Therefore, robustly segmenting images of any quality remains challenging.
@@ 2.3 Segment Anything Model and Variants
Based on SAM, some improved variants have been proposed, which can be roughly categorized into two routes. One route [medical-sam-adapter, samadapter, rsprompter, polyp-sam, vns-sam] involves using SAM for specific downstream tasks through domain-specific finetuning. These efforts typically focus on improving SAM’s performance on a specific task or dataset while sacrificing the model’s inherent generalization capabilities. Another route [samhq, robustsam, asam, mobilesam, efficientsam] is to extend SAM’s capabilities, preserving its strong generalization performance. For example, HQ-SAM [samhq] has improved SAM’s segmentation quality for objects with complex structures by adding adaptation layers while freezing SAM’s original parameters. ASAM [asam] enhances SAM’s generalization capabilities through adversarial training. Our approach follows the second path. However, diverging from these existing methods, our method focuses on improving the degradation robustness of SAM and introduces a generative latent space enhancement method within the SAM framework.
@@ 3.3.2 Channel Expansion for Head-tail Layers
We explore various methods to solve this problem (as shown in Sec. 5.3.3), including fine-tuning new head and tail layers or introducing auxiliary encoder–decoder modules to project segmentation features into a 4-dimensional space. However, these approaches proved ineffective in practice, likely due to the difficulty of aligning newly initialized layers with the pre-trained U-Net parameters while maintaining the model’s generalization capacity. To overcome this, we propose a Channel Replication and Expansion (CRE) strategy. Concretely, we replicate and concatenate the pre-trained head and tail weights of the U-Net, thereby expanding the input and output channels from 4 to 256 without altering the original parameter distribution. During training, the replicated head and tail layers are kept frozen to preserve their pre-trained generalization, while learnable LoRA [lora] layers are added to adapt to segmentation features. This design not only ensures compatibility with SAM’s latent space but also minimizes the number of learnable parameters, enabling efficient fine-tuning without sacrificing the robustness of the pre-trained generative backbone.
@@ 3.4.1 Degradation-aware Prediction Module
3.4.2 Degradation-aware Latent Space Reconstruction
The predicted $s$ functions as an explicit control signal for the generative enhancement process. In diffusion-based denoising, the parameter $\overline{\alpha}_{t}$ controls the signal-to-noise ratio at timestep $t$, as shown in Eq. 6. Our framework adaptively modulates this parameter based on $s$, effectively mapping inputs of varying quality to different positions on the diffusion trajectory. Rather than directly replacing $\overline{\alpha}_{t}$ with $s$, we perform a linear interpolation between a pre-defined minimum and maximum noise level:
@@ 3.5 Training Method
We employ a two-step fine-tuning process. In the first, we fine-tune the denoising U-Net to reconstruct high-quality features. In the second step, we fine-tune the decoder with the restored features to further align the feature space for more accurate segmentation. Detailed training and inference schemes are shown in Algorithm 1 and 2.
@@ 4.1 Multi-level Degradation Modeling
  Set Type   Category   Source data   Num.
@@
  Evaluation Set   Seen   ThinObject-5K (test set) [thin]   500
@@
  Unseen   ECSSD [ecssd]   1,000
  COCO-val [coco]   5,000
  Robust-Seg [robustsam]   1,000
@@ 4.2 Dataset Composition
Based on the above multi-level degradation model, we construct LQSeg to train our model and evaluate the segmentation performance on different levels of low-quality images. The images in LQSeg are sourced from several well-known existing datasets in the community with our synthesized degradation. The dataset details are shown in Tab. 1. In detail, for the training set, we utilize the entire training sets of LVIS [lvis], ThinObject-5K [thin], and MSRA10K [msra] as the source data and procedurally synthesize corresponding low-quality images. The evaluation set is divided into seen and unseen parts to evaluate the model’s segmentation and zero-shot generalization ability comprehensively. The seen datasets are sourced from the whole test sets of ThinObject-5K [thin] and LVIS [lvis]. The unseen parts are sourced from the entire ECSSD [ecssd] and the validation set of COCO [coco]. We generate three levels of degraded images to thoroughly assess the model’s robustness.
We further use RobustSeg [robustsam] and BDD-10K [bdd100k] datasets to evaluate the model’s performance in the presence of degradations not included in the training set. BDD-10K includes a variety of real-world degradations such as low light, blur, rain, and snow, which ensures a thorough evaluation of GleSAM++’s robustness in realistic scenarios.
@@ 5.1.2 Comparison Baselines
We compare our method with a set of comparison baselines to quantify the performance gains. Besides SAM [sam] and SAM2 [sam2], we also compare with the RobustSAM [robustsam], which has improved robustness on the degraded dataset. Additionally, we compare with two-stage methods, i.e., reconstructing images first with image reconstruction (IR) networks and passing the restored clear images to the SAM and SAM2. We use two state-of-the-art IR networks for comparison: PromptIR [promptir], and diffusion-based DiffBIR [diffbir]. Furthermore, we extended our experiments to include fine-tuning of PromptIR and DiffBIR using our degraded-clear image pairs to more comprehensively validate the superiority of our approach, referred to as PromptIR-FT and DiffBIR-FT. We employ three metrics to assess our model’s performance, including Intersection over Union (IoU), Dice Coefficient (Dice), and Pixel Accuracy (PA).
Table 2: Performance comparison on the test set of Thinobject-5K [thin] and LVIS [lvis] datasets (seen datasets) with different levels of degradation. From LQ-1 to LQ-3, the degree of degradation increases progressively. We report IoU and Dice for comparison. Our GleSAM++ and GleSAM2++ consistently outperform other competitors, especially on the most challenging LQ-3 version. The words with boldface indicate the best results, and those underlined indicate the second-best results.
@@
Table 3: Zero-shot performance comparison on the ECSSD [ecssd] and COCO [coco] datasets (unseen datasets) with different levels of degradation. These results indicate that GleSAM++/GleSAM2++ possess significant robustness in zero-shot segmentation across different levels of degradation.
@@
Table 4: Zero-shot performance comparison on Robustseg-style [robustsam] and real-world degradations. Note that our methods are not trained on such degradations. Performance is tested on the unseen ECSSD, COCO, and BDD-10K datasets. The superior performance of our method demonstrates robustness against various degradations.
@@
Figure 8: Visual comparisons on the unseen ECSSD, Robust-Seg, and BDD-10K datasets. The results demonstrate the superior generalization capability of GleSAM++ to handle unseen degradations not included in the training set.
@@ 5.2 Performance Comparisons
In this experiment, we evaluate the performance on the test set of our LQSeg, including seen-set (Tab. 2) and unseen-set (Tab. 3) evaluations.
5.2.1 Main Results on Seen and Unseen Datasets
In Tab. 2, we evaluate the performance of our GleSAM++ and GleSAM2++ on two seen datasets: ThinObject-5K [thin] and LVIS [lvis]. Each dataset contains three levels of degradation. Our GleSAM++/GleSAM2++ achieve the best performance across all three degradation levels on both datasets. Specifically, on ThinObject-5K (LQ-3), GleSAM++ improves IoU by +13% and Dice by +11% over the baseline (SAM). The performance gap widens as degradation severity increases (from LQ-1 to LQ-3), indicating that our degradation-aware generative adaptation substantially enhances model robustness. For GleSAM2++, the advantage is also evident: it surpasses SAM2 by +6% IoU and +8% Dice under the LVIS dataset (LQ-3), confirming its capability in maintaining segmentation accuracy under complex degradations. Furthermore, GleSAM++ and GleSAM2++ also outperform their predecessors, i.e., GleSAM and GleSAM2, demonstrating the improved robustness and generalization.
Tab. 3 presents the zero-shot segmentation performance of GleSAM++/GleSAM2++ on two unseen datasets: ECSSD [ecssd] and COCO [coco]. GleSAM++ and GleSAM2++ also consistently outperform other baseline methods, particularly on the most challenging LQ-3 version, underscoring their strong zero-shot generalization capabilities and potential for real-world applications.
@@ 5.2.3 Visualizations
Fig. 8 showcases the qualitative comparisons of SAM, RobustSAM, GleSAM, and GleSAM++ on the unseen dataset under various degradations. Due to the challenging degradations, previous SAM and the enhanced RobustSAM struggle to segment these objects accurately, resulting in serious detail missing and erroneous background prediction, showing their limitations. GleSAM achieves noticeably better segmentation results by introducing generative enhancement. However, its fixed-strength reconstruction limits adaptability across varying degradation severities. In contrast, GleSAM++ effectively recovers finer details and achieves more precise segmentation results. These visual results highlight the robustness of our method in processing low-quality images.
Table 5: Ablation study of each component in the proposed method, evaluated on the unseen ECSSD and COCO datasets. Each additional component positively affects the performance, demonstrating the effectiveness of the proposed methods.
@@
Table 6: Effect of Fine-tuning SAM. The performance is evaluated on the unseen ECSSD dataset. “FT-T” and “FT-D” indicate fine-tuning the SAM’s mask token and decoder, respectively. “Clear” indicates the results on the original clear images.

### v2

@@ 1 Introduction
Accurate visual perception  [maskrcnn, pcod, hybridnet, videoseg, hanet, effective, glesam] that can operate under diverse and unpredictable real-world conditions remains a fundamental task in computer vision research. Robust performance is especially critical in various high-level visual applications, such as robotics, autonomous driving, and medical image analysis. The recently developed Segment Anything Models (SAMs), including SAM [sam] and SAM2 [sam2], serving as a foundational segmentation model, have gained significant influence within the community  [rsprompter, sam-med, sam-med2, med-sam2, sa2va] due to their outstanding zero-shot segmentation abilities. It can interactively segment any object in an image using visual prompts such as points and bounding boxes. SAM’s robust generalizability has led to breakthroughs and new paradigms in various downstream tasks, including remote sensing [rsprompter, sam-rs2, sam-rs3], automatic data annotation [samrs, sam-label, p2p], and medical image segmentation [sam-med, sam-med2].
@@
A preliminary version of this work (GleSAM) was published in CVPR2025 [glesam]. This paper presents a substantial extension with the following key contributions. (1) First, we propose GleSAM++, an enhanced version with the key principle of decoupling the reconstruction process into degradation-level prediction and degradation-aware reconstruction. This design addresses a key limitation of the previous model, thereby further improving performance across any-quality images. (2) Second, we introduce a degradation-aware adaptive enhancement mechanism that dynamically regulates the reconstruction intensity in the latent space based on degradation perception, enabling a more targeted enhancement process. (3) Furthermore, we conduct a more extensive experimental evaluation across a broader range of aspects and scenarios, including robustness under diverse prompts and performance in real-world settings, thereby providing a more systematic validation of the model’s effectiveness. (4) We also provide additional analysis of the underlying factors contributing to GleSAM++’s effectiveness, along with detailed algorithmic descriptions of both the training and inference processes. (5) Finally, we present comprehensive ablation studies and visual explanations, offering deeper insights into the mechanisms behind GleSAM++.
@@ 2.1 Segmentation on Low-Quality Images
Executing robust segmentation across various scenarios is a critical issue. Numerous studies [benchmark-robustness, improve-robust, sam-robustness1, robustsam] have highlighted significant performance degradation in conventional segmentation models and foundational SAMs when confronted with low-quality images with degradation. Many related studies [improve-robust, benchmark-robustness, lwfa, fifo, dense-gram] have been proposed to enhance the robustness of segmentation models against low-quality data. For instance, URIE [urie] enhances segmentation robustness under multiple image impairments by introducing classification-based constraints. QualNet [QualNet] achieves quality-agnostic feature extraction through a reversible encoding scheme. Meanwhile, FIFO [fifo] promotes the learning of fog-resilient features in segmentation frameworks via a fog-pass filtering mechanism. These methods primarily consider a single type of degradation. Segment Anything Model (SAM) [sam] exhibits strong generalization ability due to large-scale pre-training on diverse data. However, this does not guarantee robustness under severe image degradations. Although diverse training data and augmentation improve invariance to moderate appearance variations, they cannot fully compensate for the loss of semantic cues caused by heavy noise, blur, or artifacts. Recently, RobustSAM [robustsam] is introduced to enhance the robustness of the SAM against multiple image degradations through anti-degradation feature learning. However, its performance also struggles when dealing with complex degradations. The real-world image noise is often too complex to be modeled by a single degradation [image-noise, image-noise2, image-noise3]. Therefore, robustly segmenting images of any quality remains challenging.
@@ 2.3 Segment Anything Model and Variants
Based on SAM, some improved variants have been proposed, which can be roughly categorized into two routes. One route [medical-sam-adapter, samadapter, sam-cod] involves using SAM for specific downstream tasks through domain-specific finetuning. These efforts typically focus on improving SAM’s performance on a specific task or dataset while sacrificing the model’s inherent generalization capabilities. Another route [samhq, robustsam, asam, mobilesam, vnssam] is to extend SAM’s capabilities, preserving its strong generalization performance. For example, HQ-SAM [samhq] has improved SAM’s segmentation quality for objects with complex structures by adding adaptation layers while freezing SAM’s original parameters. ASAM [asam] enhances SAM’s generalization capabilities through adversarial training. Our approach follows the second path. However, diverging from these existing methods, our method focuses on improving the degradation robustness of SAM and introduces a generative latent space enhancement method within the SAM framework.
@@ 3.3.2 Channel Expansion for Head-tail Layers
A straightforward solution is to learn new projection layers or newly initialized head/tail layers to bridge this mismatch. However, such strategies are suboptimal not only empirically (as shown in Sec. 5.3.3) but also from the perspective of compatibility with the pretrained diffusion model. The head and tail layers of the U-Net are jointly optimized with the internal denoising backbone during large-scale pretraining, and therefore form a coupled parameterization with the whole generative model. If these layers are replaced by newly initialized mappings, the transformed SAM features may exhibit activation statistics that deviate substantially from those expected by the pretrained U-Net, making the pretrained generative prior harder to transfer effectively and weakening the generalization ability of the backbone. To overcome this, we propose a Channel Replication and Expansion (CRE) strategy. Concretely, we replicate and concatenate the pre-trained head and tail weights of the U-Net, thereby expanding the input and output channels from 4 to 256 without altering the original parameter distribution. During training, the replicated head and tail layers are kept frozen to preserve their pre-trained generalization, while learnable LoRA [lora] layers are added to adapt to segmentation features. This design not only ensures compatibility with SAM’s latent space but also minimizes the number of learnable parameters, enabling efficient fine-tuning without sacrificing the robustness of the pre-trained generative backbone.
@@ 3.4.1 Degradation-aware Prediction Module
3.4.2 Degradation-aware Latent Space Reconstruction
The predicted $s$ functions as an explicit control signal for the generative enhancement process. In diffusion-based denoising, the parameter $\overline{\alpha}_{t}$ controls the signal-to-noise ratio at timestep $t$, as shown in Eq. 6. Our framework adaptively modulates this parameter based on $s$, effectively mapping inputs of varying quality to different positions on the diffusion trajectory.
@@ 3.4.2 Degradation-aware Latent Space Reconstruction
Rather than directly replacing $\overline{\alpha}_{t}$ with $s$, we perform a linear interpolation between a pre-defined minimum and maximum noise level:
@@ 3.5 Training Method
We employ a two-stage fine-tuning strategy for stable and efficient optimization. In the first stage, we adapt the denoising U-Net to restore degraded semantic features in a parameter-efficient manner, while keeping the SAM image encoder frozen. In the second stage, we fine-tune the mask decoder using the restored features for accurate segmentation. This staged design does not increase practical training complexity, since only a limited number of parameters are optimized. Instead, it improves optimization stability by decoupling feature restoration from segmentation learning. Detailed training and inference schemes are shown in Algorithm 1 and 2.
@@ 4.1 Multi-level Degradation Modeling
Set Type Evaluation Category Source data Num.
@@
Evaluation Set In-distribution ThinObject-5K (test set) [thin] 500
@@
Cross-dataset ECSSD [ecssd] 1,000
COCO-val [coco] 1,392
Out-of-distribution Robust-Seg [robustsam] 1,000
@@ 4.2 Dataset Composition
Based on the above multi-level degradation model, we construct LQSeg to train our model and evaluate the segmentation performance on different levels of low-quality images. The images in LQSeg are sourced from several well-known existing datasets in the community with our synthesized degradation. The dataset details are shown in Tab. 1. In detail, for the training set, we utilize the entire training sets of LVIS [lvis], ThinObject-5K [thin], and MSRA10K [msra] as the source data and procedurally synthesize corresponding low-quality images. The evaluation set is organized into three progressively challenging protocols to comprehensively assess both segmentation accuracy and generalization ability: 1) In-distribution evaluation. We use the test splits of ThinObject-5K [thin] and LVIS [lvis], whose source datasets are involved in training. 2) Cross-dataset evaluation. We include ECSSD [ecssd] and a de-overlapped subset of COCO-val [coco]. Since LVIS [lvis] annotations are built upon COCO images, we explicitly remove all COCO-val images that appear in the LVIS training set to eliminate any image-level data leakage. The remaining COCO-val images share no overlap with our training data. 3) Out-of-distribution evaluation. We further use RobustSeg [robustsam] and BDD-10K [bdd100k] to evaluate the model under degradation types not included in the training set. BDD-10K includes a variety of real-world degradations such as low light, blur, rain, and snow, which ensures a thorough evaluation of GleSAM++’s robustness in realistic scenarios. For each source image in protocols 1) and 2), we systematically generate three levels of degraded images to thoroughly assess the model’s robustness.
@@ 5.1.2 Comparison baselines
We compare our method with a set of comparison baselines to quantify the performance gains. 1) SAM [sam] and SAM2 [sam2], serving as the standard foundation segmentation baselines. While SAM2 is originally designed for video segmentation, our use of SAM2 here is intended to verify the compatibility of the proposed latent enhancement framework with a stronger backbone at the frame representation level. The current evaluation therefore focuses on degraded-image segmentation, and does not constitute a systematic study of temporal consistency under degraded video inputs. 2) Besides SAMs, we also compare with the RobustSAM [robustsam], which has improved robustness on the degraded dataset. 3) Additionally, we compare with two-stage methods, i.e., reconstructing images first with image reconstruction (IR) networks and passing the restored clear images to the SAM and SAM2. We use two state-of-the-art IR networks for comparison: PromptIR [promptir], and diffusion-based DiffBIR [diffbir]. 4) Furthermore, we extended our experiments to include fine-tuning of PromptIR and DiffBIR using our degraded-clear image pairs to more comprehensively validate the superiority of our approach, referred to as PromptIR-FT and DiffBIR-FT. We employ three metrics to assess our model’s performance, including Intersection over Union (IoU), Dice Coefficient (Dice), and Pixel Accuracy (PA).
Table 2: Performance comparison on the test set of Thinobject-5K [thin] and LVIS [lvis] datasets (In-distribution evaluation) with different levels of degradation. From LQ-1 to LQ-3, the degree of degradation increases progressively. We report IoU and Dice for comparison. Our GleSAM++ and GleSAM2++ consistently outperform other competitors, especially on the most challenging LQ-3 version. The words with boldface indicate the best results, and those underlined indicate the second-best results.
@@
Table 3: Performance comparison on the ECSSD [ecssd] and COCO [coco] datasets (cross-dataset evaluation) with different levels of degradation. These results indicate that GleSAM++/GleSAM2++ possess significant robustness in cross-dataset segmentation across different levels of degradation.
@@
Table 4: Zero-shot performance comparison on Robustseg-style [robustsam] and real-world degradations. Note that our methods are not trained on such degradations. Performance is tested on the ECSSD, COCO, and BDD-10K datasets. The superior performance of our method demonstrates robustness against various degradations.
@@
Figure 8: Visual comparisons on the ECSSD, Robust-Seg, and BDD-10K datasets. The results demonstrate the superior generalization capability of GleSAM++ to handle unseen degradations not included in the training set.
@@ 5.2 Performance Comparisons
In this experiment, we evaluate the performance on the test set of our LQSeg, including in-distribution (Tab. 2) and cross-dataset (Tab. 3) evaluations.
5.2.1 Main results
In Tab. 2, we evaluate the performance of our GleSAM++ and GleSAM2++ on two in-distribution datasets: ThinObject-5K [thin] and LVIS [lvis]. Each dataset contains three levels of degradation. Our GleSAM++/GleSAM2++ achieve the best performance across all three degradation levels on both datasets. Specifically, on ThinObject-5K (LQ-3), GleSAM++ improves IoU by +13% and Dice by +11% over the baseline (SAM). The performance gap widens as degradation severity increases (from LQ-1 to LQ-3), indicating that our degradation-aware generative adaptation substantially enhances model robustness. For GleSAM2++, the advantage is also evident: it surpasses SAM2 by +6% IoU and +8% Dice under the LVIS dataset (LQ-3), confirming its capability in maintaining segmentation accuracy under complex degradations. Furthermore, GleSAM++ and GleSAM2++ also outperform their predecessors, i.e., GleSAM and GleSAM2, demonstrating the improved robustness and generalization.
Tab. 3 presents the segmentation performance of GleSAM++/GleSAM2++ on two cross-dataset evaluations: ECSSD [ecssd] and COCO [coco]. GleSAM++ and GleSAM2++ also consistently outperform other baseline methods, particularly on the most challenging LQ-3 version, underscoring their strong generalization capabilities and potential for real-world applications.
@@ 5.2.2 Validation with other degradations
5.2.3 Validation of general segmentation ability
Figure 9: Comparison of SAM and GleSAM++ on clear-image benchmarks.
In this part, we further examine whether GleSAM++ preserves the general segmentation ability of SAM after domain-specific fine-tuning. The results on clear-image benchmarks are shown in Fig. 9. We conduct the evaluation on ThinObject-5K and ECSSD, and further include DUTS [duts] as an additional unseen benchmark to assess cross-dataset generalization. It can be observed that, even after domain-specific finetuning, GleSAM++ still maintains strong general segmentation performance, and consistently outperforms the original SAM across all evaluated datasets.
This observation indicates that the proposed method does not sacrifice the original general-purpose capability of the foundation model. We attribute this property to two factors. 1) GleSAM++ performs generative feature enhancement rather than replacing the original segmentation representation, and thus tends to preserve high-quality features when the input is already clear. 2) The SAM image encoder remains frozen during training, which helps retain the general visual representation learned from large-scale pre-training. As a result, the proposed fine-tuning strategy improves robustness to degraded inputs while maintaining, and in some cases even enhancing, the general segmentation ability on clear images.
@@ 5.2.4 Visualizations
Fig. 8 showcases the qualitative comparisons of SAM, RobustSAM, GleSAM, and GleSAM++ under various degradations. Due to the challenging degradations, previous SAM and the enhanced RobustSAM struggle to segment these objects accurately, resulting in serious detail missing and erroneous background prediction, showing their limitations. GleSAM achieves noticeably better segmentation results by introducing generative enhancement. However, its fixed-strength reconstruction limits adaptability across varying degradation severities. In contrast, GleSAM++ effectively recovers finer details and achieves more precise segmentation results. These visual results highlight the robustness of our method in processing low-quality images.
Table 5: Ablation study of each component in the proposed method, evaluated on the ECSSD and COCO datasets. Each additional component positively affects the performance, demonstrating the effectiveness of the proposed methods.
@@
Table 6: Effect of Fine-tuning SAM. The performance is evaluated on the ECSSD dataset. “FT-T” and “FT-D” indicate fine-tuning the SAM’s mask token and decoder, respectively. “Clear” indicates the results on the original clear images.